OpenStudy (anonymous):

could someone please help me with this question? i got 0.5.. but the back of my book says -0.5 im not sure if it's a typo, or if i did it wrong? \[\log_{2}\sqrt{36} - \log_{2} \sqrt{72}\]

jimthompson5910 (jim_thompson5910):

\[\log_{2}\sqrt{36} - \log_{2} \sqrt{72}\] \[\log_{2}\left(\frac{\sqrt{36}}{\sqrt{72}}\right)\] \[\log_{2}\left(\sqrt{\frac{36}{72}}\right)\] See where to go from here?

OpenStudy (anonymous):

equate √(36/72) ?

jimthompson5910 (jim_thompson5910):

\[\log_{2}\sqrt{36} - \log_{2} \sqrt{72}\] \[\log_{2}\left(\frac{\sqrt{36}}{\sqrt{72}}\right)\] \[\log_{2}\left(\sqrt{\frac{36}{72}}\right)\] \[\log_{2}\left(\sqrt{\frac{1}{2}}\right)\] \[\log_{2}\left(\frac{\sqrt{1}}{\sqrt{2}}\right)\] \[\log_{2}\left(\frac{1}{\sqrt{2}}\right)\] \[\log_{2}\left(\frac{1}{2^{\frac{1}{2}}}\right)\] \[\log_{2}\left(2^{-\frac{1}{2}}\right)\] How about now?

OpenStudy (anonymous):

u could convert them to base 10 and then calculate....

OpenStudy (anonymous):

uh... how do you do that? :S

jimthompson5910 (jim_thompson5910):

\[\log_{2}\sqrt{36} - \log_{2} \sqrt{72}\] \[\log_{2}\left(\frac{\sqrt{36}}{\sqrt{72}}\right)\] \[\log_{2}\left(\sqrt{\frac{36}{72}}\right)\] \[\log_{2}\left(\sqrt{\frac{1}{2}}\right)\] \[\log_{2}\left(\frac{\sqrt{1}}{\sqrt{2}}\right)\] \[\log_{2}\left(\frac{1}{\sqrt{2}}\right)\] \[\log_{2}\left(\frac{1}{2^{\frac{1}{2}}}\right)\] \[\log_{2}\left(2^{-\frac{1}{2}}\right)\] \[-\frac{1}{2}\log_{2}\left(2\right)\] \[-\frac{1}{2}(1)\] \[-\frac{1}{2}\] So \[\log_{2}\sqrt{36} - \log_{2} \sqrt{72} = -\frac{1}{2}\]

jimthompson5910 (jim_thompson5910):

I'm using the idea that \[\Large \log_{b}(X)-\log_{b}(Y)=\log_{b}\left(\frac{X}{Y}\right)\]

jimthompson5910 (jim_thompson5910):

Also, \[\Large \log_{b}(x^{y})= y\log_{b}(x)\]

OpenStudy (anonymous):

convert to base 10 as follows :